OpenStudy (iwanttogotostanford):

SIMPLE: If the original price of an item was $30.00 and Joan only paid $24.00 for it, what percentage discount did Joan receive on her purchase?

OpenStudy (iwanttogotostanford):

@Cardinal_Carlo

OpenStudy (anonymous):

So, Joan paid $6 less for what was a $30 item. We just divide the difference by the total, and then multiply by 100: \[p = \frac{ 6 }{ 30 } \times 100 = ~?\] \[p = 20\]

OpenStudy (iwanttogotostanford):

what would you do if this problem was asking for increase? Would the equation the the same then? @Cardinal_Carlo

OpenStudy (anonymous):

Yes. It would. Just remember: \[p = \frac{ change }{ total } \times 100\]

OpenStudy (iwanttogotostanford):

ok